OpenStudy (anonymous):

the square root of a neg number in terms of i -60 -169

OpenStudy (phi):

use this idea \[ \sqrt{ a \cdot b} = \sqrt{a} \sqrt{b} \] to re-write each. For example \[ \sqrt{-60} = \sqrt{-1 \cdot 60} \] now use the rule to write \[ \sqrt{-1} \sqrt{60} \] finally, use the "short name" i to replace \( \sqrt{-1} \) \[ i \sqrt{60} \] you can simplify 60 to 2*2*3*5 which has a pair of 2's that can "come out" of the radical in other words, another way to write this is \[ 2 i \sqrt{15} \]